OpenStudy (xmoses1):

The sum of the interior angles of a polygon is 540°. How many sides does the polygon have? i would like to know how to do this also

OpenStudy (anonymous):

For all polygons, the sum if the interior angles is equal to (n-2)*180 degrees where n is the number of sides. So set 540=(n-2)180 div by 180 3=n-2 add 2 n=5 The polygon has 5 sides.

OpenStudy (xmoses1):

OK that helps ALOT. Thanks for your help! :)

OpenStudy (xmoses1):

How many degrees are there in an interior angle of a regular 6-sided polygon?

OpenStudy (anonymous):

Remember (n-2)*180 so set n=6 (6-2)*180 4*180 720

OpenStudy (xmoses1):

Oh OK so i already know the formula and know that 6 is n, then i can just insert 6 for n and then work the problem out?

OpenStudy (anonymous):

Yep

OpenStudy (xmoses1):

alrighty thanks!